OpenStudy (anonymous):

what is the derivative of cos(2x)/logx

OpenStudy (psymon):

by log x do you mean natural log (lnx) or base 10?

OpenStudy (anonymous):

natural log

OpenStudy (anonymous):

nooo base

OpenStudy (anonymous):

I meant base

OpenStudy (psymon):

Well, when no base is written its an invisible 10. But a lot of people write logx to mean lnx. Ijust want to make sure which it is.

OpenStudy (anonymous):

yeah its the one with 10

OpenStudy (psymon):

Okay, cool. So we have to do quotient rule then \[\frac{ f'(x)g(x) - f(x)g'(x) }{ [g(x)]^{2} }\] Now the derivatives of the logs that are not ln work like this: \[\frac{ d }{ dx }\log _{a}u = \frac{ 1 }{ (lna)u }*\frac{ du }{ dx }\] Okay, so first we have derivative of cos(2x) which is -2sin2x, times our logx. Then minus cos(2x) timesderivative of logx, which would be 1/x(ln10). All of this is divided by (logx)^2 \[\frac{ -2\sin(2x)logx - \frac{ \cos(2x) }{ xln(10) } }{ (logx)^{2} }\] How much do you think you need to have it simplified?

OpenStudy (anonymous):

I got that :)) :)) I couldn't figure out what to do with the fraction on the top

OpenStudy (anonymous):

yes simplified please

OpenStudy (psymon):

Well, I guess we can play with it and see what happens.

OpenStudy (psymon):

\[\frac{ -2xsin(2x)\ln(10)-\cos(2x) }{ xln(10)(logx)^{2} }\] Im not even sure if theres nything worth doing to that.

OpenStudy (anonymous):

Thank you... for trying

OpenStudy (anonymous):

can I do another problem and show you if did it right?

OpenStudy (psymon):

Well, theres honestly nothing that I can imagine even trying. Unless youre sure it can be dumbed down, it just doesnt looklike it. And yeah, go for it.

OpenStudy (anonymous):

yay :D ok the problem is find y' of sin(5x)cos(5x) I got an answer which is (cos5x)5cos(5x) + 5sin(5x)-5sinx(5x)

OpenStudy (anonymous):

I have a feeling it should be simplified but it seems like the simplifying part is harder than finding the derivatives

OpenStudy (psymon):

Its supposed to beharder. Thats not just you. Calculus is often easy, its the algebra after that sucks xDD But alright, lets see \[5\cos^{2}(5x) -5\sin^{2}(5x)= 5[\cos^{2}(5x)-\sin^{2}(5x)] = 5\cos(10x)\]

OpenStudy (anonymous):

I know right! And I see what you did thanks. :D

OpenStudy (psymon):

Alright, cool xD Yeah, that one actually looked like it could be simplified, the otherone no xD

OpenStudy (anonymous):

Lol true! simplifying and deciding whether to use product rule and chain rule kills me.

OpenStudy (psymon):

well, think of it this way. You always use chain rule, its just a lot of the time the chain rule leads to multiplication by 1, which is useless.

OpenStudy (anonymous):

smh yeah it just confuses me sometimes, because I start with chain rule and then i have to do product rule at some point

OpenStudy (psymon):

If one of the layers happens to bea product rule then yeah.

OpenStudy (anonymous):

But I think I got it now I just need to learn how to simplify my work

OpenStudy (psymon):

alrighty, cool :3

OpenStudy (anonymous):

you got time for more problems?

OpenStudy (psymon):

yeah, ill see what I can do.

OpenStudy (anonymous):

cool how do you find y' of 3sin(8x)cos(8x)

OpenStudy (psymon):

Its like the same problem. I betcha the answer is 24 cos(16x), haha.

OpenStudy (anonymous):

LOL we'll see

OpenStudy (psymon):

Try to work it out and see xD

OpenStudy (anonymous):

I'm used to having the exponent so i can multiply by the 3 in this case

OpenStudy (anonymous):

I'm debatingif its going to be 0 times sin8x

OpenStudy (psymon):

Well, you can ignore the 3 until the end if you want. Its just in the way as of now, as long as you remember to multiply it in at the end.

OpenStudy (psymon):

theres no way it would be.

OpenStudy (anonymous):

oooooh im slow it should've just simply went straight to used the product rule, why im i thinking of the chain rule!

OpenStudy (psymon):

Technically thereis a chain rule everytime you take the derivative of each trig function, but....best not to think of that, just do product rule.

OpenStudy (anonymous):

im at cos^2(8x)+3sin^(8x) without the the 3 in front of cos

OpenStudy (psymon):

I guarantee you should have 8s out there somewhere xD derivative of sin(8x) = 8cos(8x) derivative of cos(8x) = -8sin(8x)

OpenStudy (anonymous):

oh yeah I wrote down the problem wrong. So I should've got 8cos^2(8x)-5sin(8x)

OpenStudy (psymon):

How a 5?

OpenStudy (anonymous):

I brought down g(X) After addition or was I suppose to ignore the 3 all the way

OpenStudy (psymon):

all the way. Multiply the 3 in at the very end.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so 8cos^2(8x)-8sin^2(8x)

OpenStudy (psymon):

Yep. You can now factor out the 8 and do like what I did above.

OpenStudy (anonymous):

im at 24(cos^2(8x)-sin^(8x) im assuming its gonna be 24cos(16x) like you said

OpenStudy (psymon):

right xDD just because it was the same thing really :P

OpenStudy (anonymous):

so for cos-sin is that the trig function and is that how you got cos

OpenStudy (psymon):

its an identity. cos^2x - sin^2x = cos2x.

OpenStudy (anonymous):

riiiight. dang i forgot my pre cal stuff :)

OpenStudy (anonymous):

thanks a millonsX you just helped me find an easy way of doing stuff I get terrified of of looking at

OpenStudy (psymon):

Yeah, you do it enough and it makes sense thankfully xD OS lagged to death, so sorry for late reply.

OpenStudy (anonymous):

no problem